Status Report

ISS On-Orbit Status 28 Jul 2002

By SpaceRef Editor
July 28, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously.  Ahead: Week 8 for Expedition Five.

Day 2 of the two-day weekend rest period.

CDR Valery Korzun took weekly readings of the SVO water supply and SP toilet flush counters in the SM for calldown to MCC-Moscow.

Today was another scheduled opportunity for the CDR to perform an attitude update of the BINS strap-down guidance system using the PUMA optical sighting device mounted in an SM window.  The two stars selected for the navigational sighting were Alpha Cassiopeiae and Beta Pegasi.  For the measurement, the station was to be put briefly in Attitude Hold from its current XPOP (x-axis perpendicular to orbit plane) orientation.

FE-2 Sergei Treschev completed the daily routine task of SOSh life support systems maintenance, while FE-1 Peggy Whitson did the "off-day" checkup of Lab payload status.

Korzun and Whitson both had their private family conferences (PFCs), via space-to-ground (S/G on S-band), with Peggy’s contact patched in by phone.

All crewmembers performed their full daily regimen of physical exercise on TVIS, RED and VELO.

Overnight, TVIS engineers at MCC-H engineers uplinked a new troubleshooting plan for the treadmill, for Korzun and Whitson to work jointly with the ground today.  The task was expected to take three hours for the two crewmembers, including a short procedure conference with the engineers before begin, if desired.

Yesterday morning at 3:00 am EDT, a bit flip in the PFCS (pump & flow control subassembly) firmware of the EETCS (early external thermal control system) loop A caused a failure warning on board, one hour after crew wake-up.  The PFCS firmware is similar to the firmware in the RPCMs (remote power control modules) and other EPS (electrical power system) boxes, so the failure response was similar. This was the first bit flip for the Loop A PFCS ORU.  It was a transient failure, and the firmware was refreshed.  The system is fully recovered.

Also yesterday, POC (Payload Operations Center) Huntsville power-cycled (turned off/on) the ARCTIC-1 payload.  This action is believed to have corrected the front panel display problem of garbled messages reported by Peggy Whitson.  At her convenience, the ARCTIC team would like her to verify that the displays are back to normal.

SpaceRef staff editor.